OpenStudy (anonymous):

If D varies directly with t and D is 108 when t is 4, find D when t is 7. 35 189 432

OpenStudy (anonymous):

D(t) = k * t D(4) = 108 = k * 4 k = 27 D(7) = 27 * 7 = 189

myininaya (myininaya):

gj ana :)